log3(x+12)-log3x=log3(x+9)
Answer by lwsshak3(11628)

log3(x+12)-log3x=log3(x+9)
**
log3(x+12)-log3(x)-log3(x+9)=0
log3(x+12)-((log3(x)+log3(x+9))=0
place under a single log
log3[(x+12)/(x)(x+9)]=0
Convert to exponential form: base(3) raised to log of number(0)=number[(x+12)/(x)(x+9)]
3^0=[(x+12)/(x)(x+9)]=1
x+12=x^2+9x
x^2+8x-12=0
let student solve for x using quadratic formula
ans:
x=-9.05 (reject, (x+9)>0)
or
x=1.29